OpenStudy (anonymous):

Line AB contains points A (0, 0) and B (2, 2). Line CD contains points C (3, 1) and D (5, 3). Lines AB and CD are parallel perpendicular neither

OpenStudy (anonymous):

@satellite73 @ganeshie8

ganeshie8 (ganeshie8):

find the slopes of both lines

ganeshie8 (ganeshie8):

if slopes are equal, then lines are parallel

OpenStudy (anonymous):

the slope for AB and for CD?

ganeshie8 (ganeshie8):

yup

OpenStudy (anonymous):

yes i got the slope of 2,2 is that correct?

ganeshie8 (ganeshie8):

A (0, 0) and B (2, 2). slope of AB = \(\large \frac{2-0}{2-0} = ?\)

ganeshie8 (ganeshie8):

C (3, 1) and D (5, 3). slope of CD = \(\large \frac{3-1}{5-3} = ?\)

ganeshie8 (ganeshie8):

are they both same ?

ganeshie8 (ganeshie8):

you meant both slopes equal to 2/2 right ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

then you're right :) since the slopes are equal, the lines will be parallel

OpenStudy (anonymous):

Thank you your the best if i was able to give you as much medal as i can i would give you alot